# TimeSeriesBench: An Industrial-Grade Benchmark for Time Series Anomaly Detection Models

### v1

@@
Abstract.
Driven by the proliferation of real-world application scenarios and scales, time series anomaly detection (TSAD) has attracted considerable scholarly and industrial interest. However, existing algorithms exhibit a gap in terms of training paradigm, online detection paradigm, and evaluation criteria when compared to the actual needs of real-world industrial systems. Firstly, current algorithms typically train a specific model for each individual time series. In a large-scale online system with tens of thousands of curves, maintaining such a multitude of models is impractical. The performance of using merely one single unified model to detect anomalies remains unknown. Secondly, most TSAD models are trained on the historical part of a time series and are tested on its future segment. In distributed systems, however, there are frequent system deployments and upgrades, with new, previously unseen time series emerging daily. The performance of testing newly incoming unseen time series on current TSAD algorithms remains unknown. Lastly, although some papers have conducted detailed surveys, the absence of an online evaluation platform prevents answering questions like “Who is the best at anomaly detection at the current stage?” In this paper, we propose TimeSeriesBench, an industrial-grade benchmark that we continuously maintain as a leaderboard. On this leaderboard, we assess the performance of existing algorithms across more than 168 evaluation settings combining different training and testing paradigms (3 current), evaluation metrics (8 current), and datasets (7 current). Through our comprehensive analysis of the results, we provide recommendations for the future design of anomaly detection algorithms. To address known issues with existing public datasets, we release an industrial dataset to the public together with TimeSeriesBench. We have also developed a comprehensive toolkit for algorithm evaluation named EasyTSAD. All code, data, and the online leaderboard have been made publicly available [^1] Code is available at https://github.com/dawnvince/EasyTSAD.
Data is available at https://github.com/CSTCloudOps/Dataset-for-TSAD.
Leaderboard is available at https://adeval.cstcloud.cn.
@@ Abstract.
copyright: acmcopyrightjournalyear: xxxxdoi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; xx; xxprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Neural networks
1. Introduction
The rapid expansion of cloud-based applications and the Web of Things (WoT) has resulted in a deluge of time-series data across a wide array of sectors, including but not limited to cloud systems (Balalaie et al., 2016; Huang et al., 2022; Dragoni et al., 2017), Web traffic (Discovery and Competition., 1999), cybersecurity (Bhatia et al., 2020), and healthcare (Moody and Mark, 2001a; Bachlin et al., 2009). This vast and complex data has brought to the fore the critical role of time series anomaly detection (TSAD), which aims to identify irregular patterns in time series data. These outliers can often signify critical deviations from the norm, such as impending machine failures in industrial processes (Su et al., 2019), malignant network traffic attacks in web applications (Ring et al., 2019), or the early signs of disease outbreaks in public health surveillance (Moody and Mark, 2001b).
Owing to the clear significance and applicability of TSAD in these real-world scenarios, in recent years, a variety of anomaly detection methods, particularly those based on deep learning, are burgeoning incessantly (Xu et al., 2022, 2018; Ahmad et al., 2017a; Chen et al., 2020; Hundman et al., 2018; Li et al., 2018; Si et al., 2023). Nonetheless, there is a considerable divergence across the results reported by different papers even on the same dataset, as they employed various evaluation criteria and learning schemas. Despite the existence of some benchmarks for anomaly detection algorithms (Emmott et al., 2013; Tatbul et al., 2018; Lai et al., 2021; Jacob et al., 2020; Paparrizos et al., 2022b; Schmidl et al., 2022), they struggle to provide practical guidance for industrial domain experts to apply and develop deep learning methods as they pay more attention to statistical methods. In general, applying TSAD algorithms to practical systems faces these main obstacles: (I) Training specific model for each time series and deploying an exclusive model for each time series results in unaffordable maintenance costs. (II) New systems are more likely to encounter issues when first launched or upgraded, which makes anomaly detection even more essential despite their almost non-existent volume of historical data. (III) Existing metrics hold one-sided assumptions for evaluating how well the algorithms perform, which cannot offer an effective reference for industrial practice. (IV) There is no platform continuously integrating new methods and comparing them in a unified and intuitive manner, similar to the GLUE Leaderboard in NLP (Wang et al., 2019), which prevents experts in the industry from keeping pace with the latest algorithms and advancements in academia.
@@ 1. Introduction
  • For the first time, we evaluate several well-known state-of-the-art (SOTA) anomaly detection methods on TimeSeriesBench and discovered some enlightening conclusions, providing directions for future optimization.
@@
  • To address the issue of inaccurate anomaly labeling in existing public datasets (Wu and Keogh, 2021), we collaborated with a global large company and invited business system experts to meticulously annotate anomalies in the online system. After detailed calibration, we have also made the dataset publicly available to the community as part of TimeSeriesBench, supplementing existing anomaly detection datasets.
Figure 1. The industrial-grade concerns and TimeSeriesBench’s solutions on benchmarking univariate time series anomaly detection algorithms.
2. PRELIMINARIES
2.1. Time Series Anomaly Detection
@@
2.2. Anomaly Types
Following the behavior-driven taxonomy (Lai et al., 2021), anomaly types can be roughly divided into point-wise outliers and pattern-wise outliers (Fig. 2). Point-wise outliers denote unexpected incidents like spikes or glitches on individual time points or within very short periods of time. Pattern-wise outliers represent anomalous subsequences that span over a certain period of time, often characterized as discordances or inconsistencies within the data.
Figure 2. Illustrations of anomaly types. Anomalous segments are highlighted in pink.
3. TimeSeriesBench Settings
@@ 3.1. Datasets
High-quality datasets are the prerequisite for effective training and reasonable performance evaluations for models. Unfortunately, several publicly available time series datasets are claimed flawed due to their unrealistic anomaly density, or mislabeled ground truth (Wu and Keogh, 2021). Moreover, the proportion of different types of outliers can vary widely among datasets as the time series are gathered from irrelevant domains and applications. For example, Yahoo (Research, 2015) is dominated by global and contextual point-wise outliers on the basis of the behavior-driven taxonomy proposed by  (Lai et al., 2021) (the trivial first-order difference can achieve a comparable performance), while pattern-wise outlier plays a significant role in UCR archive (Wu, 2022a) (shown in Sec. 4). To this end, we gather multiple well-labeled real-world datasets covering many application domains to diversify the distributions of anomalies, meanwhile excluding markedly flawed datasets (some cases in  (Wu and Keogh, 2021)). Furthermore, we introduce a synthetic dataset generated from TODS (Lai et al., 2021) for the convenience of specific case analysis due to its good interpretability, because the time series adheres to well-designed distributions.
@@
Real-world datasets. We employ AIOPS (Competition, 2018), WSD (Zhang et al., 2022a), Yahoo (Research, 2015), NAB (Ahmad et al., 2017b), and UCR (Wu, 2022a) for evaluations. Real-world datasets are more susceptible to uncertainties and tend to mix up with even inexplicable noise, which raises the demand in terms of model robustness. As the classes of anomalies are imbalanced among these datasets, practitioners focusing on the general-purpose model should refer to detection performance on all datasets, while application-oriented tasks need to pay more attention to performance on datasets that are strikingly similar to specific scenes. Meanwhile, we release a new dataset collected from the production environment, called NEK (Network Equipment KPI), for a more comprehensive evaluation.
@@
Synthetic dataset. TODS (Lai et al., 2021) introduces a novel taxonomy, categorizes anomalies into five types, and publishes an anomaly generation toolkit in line with their claims. We conduct modifications based on their codes to generate longer and non-trivial anomalies, which would allow researchers to determine if models are capable of particular types of anomalies in a straightforward manner.
3.2. Learning Schema
@@
Existing benchmarks take it for granted that the training process of the detector should comply with the naive task-specific schema, more specifically, training an exclusive detector for each time series solely leveraging its own historical patterns. We claim that it’s time to break the prevailing stereotype, especially for deep learning models, given the great transformations observed in application scenarios during recent years. On the one hand, with the maturation and improvement of surveillance systems in various fields, the quantity of time series to be monitored has experienced a substantial increase, reaching several orders of magnitude higher than before. This significantly raised the storage overheads if detectors are exclusive. On the other hand, the previous years witnessed many seminal works pertinent to foundational models for time series forecasting or other general tempo tasks (Anonymous, 2024b, c; Garza and Mergenthaler-Canseco, 2023; Zhou et al., 2023). We aspire to establish a foundation for the standardized training and evaluation of large-scale, general-purpose time series models, especially those designed for anomaly detection. Thus we introduce two novel learning schemas, called all-in-one mode and zero-shot mode, to assess the performance variations of models in scenarios involving large-scale data or zero-shot settings.
@@
All-in-one schema. In this mode, only one unified model instance is trained using all the sequences in the dataset, and the trained model is then applied for real-time anomaly detection across all sequences within the dataset. This schema exposes more patterns embedded in various series to the model, thereby providing an additional opportunity for the model to learn common and inherent traits shared among the time series. However, due to the differing definitions of anomalies across series, this carries the risk of confounding the model with conflicting information when detecting anomalies online (i.e., an anomaly present in a specific curve may not necessarily be considered an anomaly in other curves). Several novel methods (Zhang et al., 2022b; Ren et al., 2019) have adopted this schema driven by practical needs.
@@
Figure 3. Illustrations of three learning schemas. $s_{n}$ indicates the n-th time series in the dataset.
3.3. Evaluation Criteria
A flawless evaluation criteria serves as the foundation for not only assessing the effectiveness of methods but also guiding model parameters optimization. Nevertheless, recent research has uncovered substantial limitations in commonly employed evaluation criteria, on which some seemingly absurd methods (e.g. noise generated by Random Guess) can outperform all others (Doshi et al., 2022; Sehili and Zhang, 2023; Kim et al., 2022; Garg et al., 2021). Also, some assumptions employed by certain evaluation criteria also conflict with the principles of industrial practice. Therefore, it is imperative to revise an anti-cheating and industrial-oriented evaluation protocol for our benchmark. It is crucial to emphasize that the selection of protocol is closely tied to the specific application context and determines how models perform with regard to the aspect you are interested in. In this paper, we focus on a generic real-time anomaly detection task.
In general, evaluation criteria employed in existing benchmarks can be roughly categorized into point-based protocols and range-based protocols. Point-based protocols like traditional F1 treat each individual data point as a separate sample, disregarding the holistic characteristics of the anomaly segment. Range-based protocols incorporate segment-level features, for instance, the detection latency (Jacob et al., 2020), into the evaluation.
In comparison to the aforementioned criteria, we prefer a protocol that aligns better with the typical workflow of real-time anomaly detection systems deployed in real-world scenarios. In practical application scenarios, whenever the anomaly score exceeds a predefined threshold, the system triggers an alert to notify the operators. Therefore, an operator tends to prioritize the following key issues:
@@ 3.3. Evaluation Criteria
To address issue a, we utilize a strategy of ”adjusting” the output of the algorithm, namely point-adjustment (PA), which is widely adopted (Qin et al., 2023; Shen et al., 2020; Su et al., 2019; Audibert et al., 2020; Li et al., 2021; Zhang et al., 2022b; Si et al., 2023). Under this strategy, all timestamps within an anomalous segment are assigned the highest anomaly score present within that segment, thus the whole anomaly segment is considered to be detected if at least one anomaly score surpasses the threshold. Then the F1 score is obtained in a point-based manner (point-wise PA in Fig. 4).
@@
However, point-wise PA introduces biased true positives and false negatives due to the point-based criteria. As shown in Fig. 4, as a long anomaly segment is detected under point-adjustment, even if two false alarms are triggered, the Precision is up to 0.8, which contradicts the demand corresponding to issue b. The intrinsic flaw of this criteria is that the detection is rewarded generously while pulsing false alarm is penalized just once (Doshi et al., 2022). Garg et al. (Garg et al., 2021) revise the primitive PA from an event perspective. Each anomaly segment is treated as an individual event and contributes to a true positive or false negative only once. This protocol, namely event-wise PA, prevents the inflated evaluation scores originated from illogical criteria, while absolutely overlooking the length of the anomaly segments. Based on the observation c , we apply a severity coefficient $\ln{(k+e)}$ to adjust the true positive and false positive measurements associated with the given anomalous segment of length $k$. The protocol is called reduced-length PA, and we categorize it along with event-wise PA as event-based evaluation criteria.
@@
Figure 4. Illustrations of evaluation protocols based on point-adjustment (PA). Point-wise PA gives an inflated score when some anomaly segments persist for a long duration. Event-wise PA treats each anomaly segment as an event, completely disregarding the length of the anomaly segment. Reduced-length PA considers the trade-offs between the two methods, holding greater practical significance in real-world applications.
Issue d demands exceptionally high data set quality, whereas some datasets fail to offer labels without positional bias. Due to the limitations of labels, we present this portion as supplementary content. More details can be found in Appendix A.
@@
Besides, most statistical and deep learning methods generate anomaly scores under prediction-based or reconstruction-based frameworks, both heavily relying on the pattern provided by previous time windows. Hence, the recently concluded anomaly segment, particularly those with longer durations, has the potential to heavily interfere with the subsequent detection process. As shown in Fig. 5, from a qualitative perspective, we observe that the method accurately detects the frequency anomaly event. However, some unexpected false positives emerge due to the aforementioned reasons, resulting in an underestimated evaluation. We slightly extended the anomaly segments (less than 10 time points) to tolerate such occurrences, to make the evaluation more in line with our intuitive comprehension. We carefully handle edge cases to avoid merging two anomaly segments during the operation.
@@
Figure 5. Illustrations of prolonging the anomaly segments. The blue line denotes the anomaly scores provided by a particular method. Considering the interference caused by anomalies on the detection of normal values, we should exclude false positives that occur immediately after an anomaly segment.
3.4. Algorithms

### v3

@@
Abstract
Time series anomaly detection (TSAD) has gained significant attention due to its real-world applications to improve the stability of modern software systems. However, there is no effective way to verify whether they can meet the requirements for real-world deployment. Firstly, current algorithms typically train a specific model for each time series. Maintaining such many models is impractical in a large-scale system with tens of thousands of curves. The performance of using merely one unified model to detect anomalies remains unknown. Secondly, most TSAD models are trained on the historical part of a time series and are tested on its future segment. In distributed systems, however, there are frequent system deployments and upgrades, with new, previously unseen time series emerging daily. The performance of testing newly incoming unseen time series on current TSAD algorithms remains unknown. Lastly, the assumptions of the evaluation metrics in existing benchmarks are far from practical demands. To solve the above-mentioned problems, we propose an industrial-grade benchmark TimeSeriesBench. We assess the performance of existing algorithms across more than 168 evaluation settings and provide comprehensive analysis for the future design of anomaly detection algorithms. An industrial dataset is also released along with TimeSeriesBench.
Index Terms:
@@ Abstract
I Introduction
In recent years, we have witnessed the rapid growth in the scale of software systems and services. To ensure service quality assurance and maintain user satisfaction, IT operation engineers must monitor the time series metrics like response time or success rate to judge if there are system failures. Time series anomaly detection (TSAD) aims to identify irregular patterns in these time series data, which indicate potential or existing system failures. The outliers can often signify critical deviations from the norm, such as impending machine failures in industrial processes [1] and malignant network traffic attacks in web applications [2]. Accurate and timely anomaly detection can help to reduce the mean time to repair, reduce the loss in revenue and maintain the reputation and branding for a company.
Owing to the clear significance and applicability of TSAD in these real-world scenarios, in recent years, a variety of anomaly detection methods, particularly those based on deep learning, are burgeoning incessantly [3, 4, 5, 6, 7, 8, 9, 10]. Nonetheless, there is a considerable divergence across the results reported by different papers even on the same dataset, as they employed various evaluation criteria and learning schemas. Despite the existence of some benchmarks for anomaly detection algorithms [11, 12, 13, 14, 15, 16], they struggle to provide practical guidance for industrial domain experts to apply and develop deep learning methods as they pay more attention to statistical methods. In general, applying TSAD algorithms to practical systems faces these main obstacles: (I) Training specific model for each time series and deploying an exclusive model for each time series results in unaffordable maintenance costs. (II) New systems are more likely to encounter issues when first launched or upgraded, which makes anomaly detection even more essential despite their almost non-existent volume of historical data. (III) Existing metrics hold one-sided assumptions for evaluating how well the algorithms perform, which cannot offer an effective reference for industrial practice. (IV) There is no platform continuously integrating new methods and comparing them in a unified and intuitive manner, similar to the GLUE Leaderboard in NLP [17], which prevents experts in the industry from keeping pace with the latest algorithms and advancements in academia.
Figure 1: The industrial-grade concerns and TimeSeriesBench’s solutions on benchmarking univariate time series anomaly detection algorithms.
@@ I Introduction
  • For the first time, we employ the cost-friendly all-in-one and zero-shot settings to evaluate several well-known state-of-the-art (SOTA) anomaly detection methods on TimeSeriesBench and discover some enlightening conclusions, providing directions for future optimization.
@@
  • To address the issue of inaccurate anomaly labeling in existing public datasets [18], we collaborated with a global large company and invited business system experts to meticulously annotate anomalies in the online system. After detailed calibration, we have also made the dataset publicly available to the community as part of TimeSeriesBench, supplementing existing anomaly detection datasets. The codes, data and the online leaderboard are released publicly [^1] The code is available at https://github.com/CSTCloudOps/EasyTSAD. The new proposed dataset is available at https://github.com/CSTCloudOps/Dataset-for-TSAD, and the online leaderboard is available at https://adeval.cstcloud.cn..
II PRELIMINARIES
II-A Time Series Anomaly Detection
@@
II-B Anomaly Types
Following the behavior-driven taxonomy [13], anomaly types can be roughly divided into point-wise outliers and pattern-wise outliers (Fig. 2). Point-wise outliers denote unexpected incidents like spikes or glitches on individual time points or within very short periods of time. Pattern-wise outliers represent anomalous subsequences that span over a certain period of time, often characterized as discordances or inconsistencies within the data.
Figure 2: Illustrations of anomaly types. Anomalous segments are highlighted in pink.
III TimeSeriesBench Settings
@@ III-A Datasets
High-quality datasets are the prerequisite for effective training and reasonable performance evaluations for models. Unfortunately, several publicly available time series datasets are claimed flawed due to their unrealistic anomaly density, or mislabeled ground truth [18]. Moreover, the proportion of different types of outliers can vary widely among datasets as the time series are gathered from irrelevant domains and applications. For example, Yahoo [19] is dominated by global and contextual point-wise outliers on the basis of the behavior-driven taxonomy proposed by  [13] (the trivial first-order difference can achieve a comparable performance), while pattern-wise outlier plays a significant role in UCR archive [20] (shown in Sec. V). To this end, we gather multiple well-labeled real-world datasets covering many application domains to diversify the distributions of anomalies, meanwhile excluding markedly flawed datasets (some cases in  [18]). Furthermore, we introduce a synthetic dataset generated from TODS [13] for the convenience of specific case analysis due to its good interpretability, because the time series adheres to well-designed distributions.
@@
Real-world datasets. We employ AIOPS [21], WSD [22], Yahoo [19], NAB [23], and UCR [20] for evaluations. Real-world datasets are more susceptible to uncertainties and tend to mix up with even inexplicable noise, which raises the demand in terms of model robustness. As the classes of anomalies are imbalanced among these datasets, practitioners focusing on the general-purpose model should refer to detection performance on all datasets, while application-oriented tasks need to pay more attention to performance on datasets that are strikingly similar to specific scenes. Meanwhile, we release a new dataset collected from the production environment, called NEK (Network Equipment KPI), for a more comprehensive evaluation.
@@
Synthetic dataset. TODS [13] introduces a novel taxonomy, categorizes anomalies into five types, and publishes an anomaly generation toolkit in line with their claims. We conduct modifications based on their codes to generate longer and non-trivial anomalies, which would allow researchers to determine if models are capable of particular types of anomalies in a straightforward manner.
III-B Learning Schema
@@
Existing benchmarks take it for granted that the training process of the detector should comply with the naive task-specific schema, more specifically, training an exclusive detector for each time series solely leveraging its own historical patterns. We claim that it’s time to break the prevailing stereotype, especially for deep learning models, given the great transformations observed in application scenarios during recent years. On the one hand, with the maturation and improvement of surveillance systems in various fields, the quantity of time series to be monitored has experienced a substantial increase, reaching several orders of magnitude higher than before. This significantly raised the storage overheads if detectors are exclusive. On the other hand, the previous years witnessed many seminal works pertinent to foundational models for time series forecasting or other general tempo tasks [24, 25, 26, 27]. We aspire to establish a foundation for the standardized training and evaluation of large-scale, general-purpose time series models, especially those designed for anomaly detection. Thus we introduce two novel learning schemas, called all-in-one mode and zero-shot mode, to assess the performance variations of models in scenarios involving large-scale data or zero-shot settings.
@@
All-in-one schema. In this mode, only one unified model instance is trained using all the sequences in the dataset, and the trained model is then applied for real-time anomaly detection across all sequences within the dataset. This schema exposes more patterns embedded in various series to the model, thereby providing an additional opportunity for the model to learn common and inherent traits shared among the time series. However, due to the differing definitions of anomalies across series, this carries the risk of confounding the model with conflicting information when detecting anomalies online (i.e., an anomaly present in a specific curve may not necessarily be considered an anomaly in other curves). Several novel methods [28, 29] have adopted this schema driven by practical needs.
@@
Figure 3: Illustrations of three learning schemas. $s_{n}$ indicates the n-th time series in the dataset.
III-C Evaluation Criteria
A flawless evaluation criteria serves as the foundation for not only assessing the effectiveness of methods but also guiding model parameters optimization. Nevertheless, recent research has uncovered substantial limitations in commonly employed evaluation criteria, on which some seemingly absurd methods (e.g. noise generated by Random Guess) can outperform all others [30, 31, 32, 33]. Also, some assumptions employed by certain evaluation criteria also conflict with the principles of industrial practice. Therefore, it is imperative to revise an anti-cheating and industrial-oriented evaluation criterion for our benchmark. It is crucial to emphasize that the selection of criterion is closely tied to the specific application context and determines how models perform with regard to the aspect you are interested in. In this paper, we focus on a generic real-time anomaly detection task.
In industrial practice, we prefer a criterion that aligns better with the typical workflow of real-time anomaly detection systems deployed in real-world scenarios. In a practical software system, whenever the anomaly score exceeds a predefined threshold, the system triggers an alert to notify the operators. Therefore, an operator tends to prioritize the following key issues:
@@ III-C Evaluation Criteria
Unfortunately, current evaluation criteria [14, 34, 30, 32, 35, 36] fail to simultaneously address all of the above issues. Thus we propose a new criterion that eliminate the effect of the factors that distort the evaluation results by the following strategies:
@@
Point Adjustment. We utilize a strategy of ”adjusting” the output of the algorithm, namely point-adjustment (PA), as the solution to issue a. This strategy is widely adopted [37, 38, 1, 39, 40, 28, 9]. Under this strategy, all timestamps within an anomalous segment are assigned the highest anomaly score present within that segment, thus the whole anomaly segment is considered to be detected if at least one anomaly score surpasses the threshold. Then the F1 score is obtained in a point-based manner (point-wise PA in Fig. 4).
@@
Anomaly Weights Revision. However, the naive point-wise PA introduces biased true positives and false negatives. As shown in Fig. 4, as a long anomaly segment is detected under point-adjustment, even if two false alarms are triggered, the Precision is up to 0.8, which contradicts the demand corresponding to issue b. The intrinsic flaw of this criteria is that the detection is rewarded generously while pulsing false alarm is penalized just once [30]. Garg et al. [33] revise the primitive PA from an event perspective. Each anomaly segment is treated as an individual event and contributes to a true positive or false negative only once. This criterion, namely event-wise PA, prevents the inflated evaluation scores originated from illogical criteria, while absolutely overlooking the length of the anomaly segments. Based on the observation c , we apply a severity coefficient $\ln{(k+e)}$ to adjust the true positive and false positive measurements associated with the given anomalous segment of length $k$. The criterion is called reduced-length PA, and we categorize it along with event-wise PA as event-based evaluation criteria.
@@
Figure 4: Illustrations of evaluation criteria based on point-adjustment (PA). Point-wise PA gives an inflated score when some anomaly segments persist for a long duration. Event-wise PA treats each anomaly segment as an event, completely disregarding the length of the anomaly segment. Reduced-length PA considers the trade-offs between the two methods, holding greater practical significance in real-world applications.
Latency Constraint. Issue d is addressed by imposing strict constraints on the detection latency. As depicted in the illustration (Fig. 5), assuming the latency limit ($k$) is set to 3, an anomaly is considered effectively detected only if it is identified within three sampling points after its occurrence. We designate this strategy as k-delay adjustment. This measure enables a more precise assessment of whether the model can meet the requirement of the scenario where there is a high demand for real-time responsiveness. It is equally essential to acknowledge that this approach is applicable only to datasets whose anomalies are labeled without positional bias. Unfortunately, not all datasets can meet this requirement, thus we present the related results as supplementary content.
Figure 5: Illustrations of k-delay adjustment. This strategy can be combined with point-adjustment as a complementary evaluation paradigm.
@@
Anomaly Lag Elimination. Most statistical and deep learning methods generate anomaly scores under prediction-based or reconstruction-based frameworks, both heavily relying on the pattern provided by previous time windows. Hence, the recently concluded anomaly segment, particularly those with longer durations, has the potential to heavily interfere with the subsequent detection process. As shown in Fig. 6, from a qualitative perspective, we observe that the method accurately detects the frequency anomaly event. However, some unexpected false positives emerge due to the aforementioned reasons, resulting in an underestimated evaluation. We slightly extended the anomaly segments (less than 10 time points) to tolerate such occurrences, to make the evaluation more in line with our intuitive comprehension. We carefully handle edge cases to avoid merging two anomaly segments during the operation.
@@
Figure 6: Illustrations of prolonging the anomaly segments. The blue line denotes the anomaly scores provided by a particular method. Considering the interference caused by anomalies on the detection of normal values, we should exclude false positives that occur immediately after an anomaly segment.
III-D Algorithms